In a family there are two cars. The sum of the average miles per gallon obtained by the two cars in a particular week is 50 . The first car has consumed 25 gallons during that week, and the second has consumed 30 gallons, for a total of 1400 miles driven by the two cars combined. What was the average gas mileage obtained by each of the two cars in that week?
Answer by richwmiller(17219)

let a =mpg one car and b= mpg the other
(a+b)=50
b=50-a
25a+30(50-a)=1400